OpenStudy (anonymous):

please help. -3/4q=12 q=?

OpenStudy (anonymous):

What well you know the equation \[\frac{ -3 }{ 4q }=12\] So you have to isolate the variable

OpenStudy (anonymous):

To do this you must multiply both sides by 4q in order to get the q out from under the fraction. Now you have \[-3=12*4q\]

OpenStudy (anonymous):

Then you can simply and multiply 12 by 4q to get \[-3=48q\] Now can you tell me what you think the answer is?

OpenStudy (anonymous):

-3 - q=12 q = ? -4

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Ohhhhh okay so let me confirm, its suppose to look like this ? \[\frac{ -3 }{ -4 }q=12\]

OpenStudy (anonymous):

ya the whole fraction is negative

OpenStudy (anonymous):

oh okay got you so let's start this over again lol So you have the equation \[-\frac{ 3 }{ 4 }q=12\] and you need to isolate the variable. All you have to do to find it is multiply 12 by the inverse of \[-\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

which should be \[-\frac{ 4}{ 3 }\] So you multiply 12 by \[-\frac{ 4}{ 3 }\] and that is the answer for q. What did you get?

OpenStudy (anonymous):

q = -16

OpenStudy (anonymous):

there ya go^ q should equal -16

OpenStudy (anonymous):

ya thats what i got thank you tbeneka

OpenStudy (anonymous):

Good job! and no problem, anytime